OpenStudy (anonymous):

Two number cubes are rolled --one blue and one yellow. Find the probability that the blue cube shows a 4 and the product is less than 20. Enter your answer in simplified fraction form; example: 3/5.

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (fibonaccichick666):

your thoughts?

OpenStudy (anonymous):

1/6

OpenStudy (fibonaccichick666):

full explanation of each step please

OpenStudy (anonymous):

then 4/6

OpenStudy (anonymous):

1/6= roll a 4

OpenStudy (fibonaccichick666):

from the start

OpenStudy (fibonaccichick666):

k good

OpenStudy (anonymous):

4/6= the amount of number you can multiply by 4 before getting 20

OpenStudy (anonymous):

4/36

OpenStudy (anonymous):

divide by 4=1/9

OpenStudy (fibonaccichick666):

continue

OpenStudy (fibonaccichick666):

and good